# Dissecting Logical Reasoning in LLMs: A Fine-Grained Evaluation and Supervision Study

### v1

@@ 6 Conclusion
Our work has two main limitations. First, although FineLogic evaluates models across three complementary dimensions, it is built upon a limited set of datasets. While we selected benchmarks that span multiple domains and reasoning depths, no fixed dataset collection can offer a complete assessment of LLM reasoning capabilities. Nevertheless, the modular design of FineLogic allows it to be easily extended to future benchmarks, enabling the same set of evaluation tasks—overall accuracy, stepwise soundness, and representation-level probing—to be applied to new problem settings. Second, although we provide detailed analyses of how different supervision styles succeed or fail across tasks, we do not explore how to systematically integrate these insights into the design of more comprehensive or hybrid SFT datasets. We leave it to future work to develop adaptive or mixed-format training strategies that balance generalization, structural soundness, and interpretability in logical reasoning.
@@ A.1 Detailed Dataset Information
In this section, we present the logical reasoning datasets used in our experiments. All datasets are publicly available. We describe the data sources and sampling procedures in detail below.
@@
(Morishita et al., 2024). The FLD dataset is constructed based on the number of reasoning steps (denoted as steps). We first generate 50,000 raw samples uniformly distributed across step depths from 0 to 15. For each step, we randomly sample 250 PROVED and 250 DISPROVED examples, yielding a total of $16\times 500=8{,}000$ labeled samples. Additionally, we include 1,500 randomly drawn UNKNOWN instances, resulting in a total training set of 9,500 samples.
The test set follows a similar sampling process but spans a broader range of reasoning depths (0–19). We select 25 PROVED and 25 DISPROVED instances per step, yielding $20\times 50=1{,}000$ labeled examples. We also sample 100 UNKNOWN instances, producing a test set of 1,100 samples.
@@
(Han et al., 2022). The FOLIO dataset contains 203 test examples, all of which are included in our experiments.
@@
(Patel et al., 2024). The Multi-LogiEval dataset consists of five subsets with different depth (denoted ‘d1’ through ‘d5’). We include only the ‘d4’ and ‘d5’ subsets to maintain a consistent dataset size, yielding a total of 390 examples. From each, we extract all samples from the First-Order Logic and Propositional Logic categories, masking their logic type annotations to ensure a uniform setting.
@@
(Saparov and He, 2022). We use the standard ProntoQA dataset rather than the OOD variant, following prior work (Pan et al., 2023). We use 500 hardest samples requiring 5-hop reasoning steps as a held-out test set. For training, we sample 3,200 problems that each require 3-hop reasoning, selected to provide sufficient multi-step depth while remaining tractable for supervised fine-tuning.
@@ A.2 Detailed Baseline Methods
(Feng et al., 2023) is a novel language model designed to emulate the reasoning processes of logical solvers. Unlike traditional solver-augmented models that parse natural language into symbolic representations before reasoning, LOGIPT directly learns to generate symbolic reasoning steps, bypassing potential parsing errors.
@@
(Creswell et al., 2022) enhances LLMs by n-shot learning between selecting relevant facts and inferring new information, enabling interpretable multi-step logical reasoning. In the experiments, we achieved rather low accuracy for limited questions (hypothesis) to trigger new inferences.
@@
(Pan et al., 2023) is a neuro-symbolic framework that addresses LLMs’ limitations in complex logical reasoning by integrating symbolic solvers with language models. The approach works through a three-stage pipeline: (1) using LLMs to translate natural language problems into symbolic formulations, (2) employing deterministic symbolic solvers to perform faithful logical inference, and (3) interpreting results back to natural language. It features a self-refinement module that iteratively revises symbolic representations based on solver error messages.
@@
(Xu et al., 2024) introduces a novel framework that enhances LLMs by integrating symbolic expressions and logical rules into the CoT reasoning process. It translates natural language contexts into symbolic formats, derives step-by-step plans using symbolic logic, and employs a verifier to ensure the correctness of translations and reasoning chains.
@@ B.3 Task 1: Correctness Spanning Steps
  • For each test sample, we identify the smallest $i$ such that the classifier correctly predicts the label at step $i$ but fails at step $i-1$.
The correctness spanning length is defined as $n-i$, capturing how early the model “knows” the correct answer.
@@ B.5 Task 3: Next-Step Derivability
Appendix C Experiment Details
This section provides further details on our experimental results. Table 5 presents a comprehensive breakdown of FLD accuracy across different reasoning step ranges for the evaluated models and settings. The data illustrates that while models fine-tuned with natural language supervision (e.g., Llama-3.1-SFT-NL achieving 89.5% accuracy for 0-3 steps on FLD) perform strongly on tasks with shallower reasoning depths, their symbolic reasoning counterparts tend to exhibit greater resilience as the complexity and number of reasoning steps increase. For instance, on FLD problems requiring 16-19 steps, Llama-3.1-SFT-Symb-Filter (62.5%) and Llama-3.1-SFT-Symb-Struct (58.5%) maintain higher accuracy compared to Llama-3.1-SFT-NL (46.0%), highlighting the benefit of symbolic formats for robust multi-step inference.
@@ Appendix C Experiment Details
Direct 40.5 30.0 24.0 27.0 25.5
CoT 41.5 32.0 29.0 24.0 19.5
Few-shot 49.5 45.5 33.0 39.0 32.0
Logic-LM 56.4 41.3 32.6 28.8 26.2
SymbCoT 57.8 41.0 39.0 37.8 35.4
Sel-Inf 63.5 55.5 52.5 45.0 42.0
LogiPT 72.5 53.5 51.0 35.0 37.0
SFT-NL 89.5 72.5 52.0 56.5 46.0
SFT-Symb-Struct 88.5 78.5 65.0 66.0 58.5
SFT-Symb-Filter 72.0 73.0 67.5 72.0 62.5
Llama-3.1 -8B-Instruct SFT-Symb-Direct 81.0 58.5 48.5 38.5 27.5
@@ Appendix F Prompt Template
2. For each step: First, write a brief, natural language explanation on its own line (e.g. "Assume for contradiction: [formula]" or "From [inputs], we derive:"). On the next line, write the step label and the logical formula as in the original proof (e.g. assump1: A(b), int2: $\neg$¬B(b), etc.). Do not put both the explanation and the formula on the same line. For assumptions, use "Assume for contradiction: [formula]" then write assumpX: [formula] on the following line. For a standard derived step, use "From [inputs], we derive:" then on the following line write intX: [formula]. For contradictions, use "Contradiction:" then on the following line write $\bot$". For reductio ad absurdum, use "By reductio ad absurdum from [step number]:" then write the derived conclusion on the next line. Do not skip formula labels or step names. Write both the explanation and the labeled formula.
@@ Conclusion:
Figure 14: Prompt template for evaluating step atomicity. Placeholders: {premises_str} (a string listing the premises), {concl_text_full} (a string representing the conclusion). The model is expected to return ‘true’ or ‘false’ indicating if the inference from premises to conclusion is a single, indivisible logical step.

### v2

@@ 7 Conclusion
Our work has two primary limitations. First, our evaluation is conducted on a fixed suite of datasets. While these were chosen for their diversity, this necessarily scopes our findings to these specific benchmarks, and the observed model behaviors may not generalize to the entire domain of logical reasoning tasks. Second, our analysis of supervision styles treats each format in isolation. Consequently, the potential benefits of hybrid training strategies, which could combine the complementary strengths of natural language and symbolic formats, remain unexplored in this study.
Acknowledgment
We are grateful to Xiuying Chen of MBZUAI for her constructive comments and helpful suggestions, which helped improve our manuscript.
@@ A.1 Detailed Dataset Information
Morishita et al. (2024) This is a synthetic dataset designed to test generalization across varying reasoning depths. Our test set consists of 1,100 samples spanning 20 reasoning depths (0–19). It is specifically constructed to include out-of-distribution problems (depths 16–19) that are more difficult than those used for fine-tuning.
@@
Han et al. (2022) A human-curated benchmark for first-order logic reasoning in natural language. We use the complete test set of 203 problems for our evaluation.
@@
Patel et al. (2024) This dataset systematically combines multiple inference rules to create complex problems. To focus on challenging multi-step reasoning, we select 390 problems from its two most difficult subsets (depths 4 and 5), using only the First-Order and Propositional Logic categories.
@@
Saparov and He (2022) This dataset evaluates multi-hop reasoning over a semi-synthetic knowledge base. For our test set, we follow the challenging setup from prior work Pan et al. (2023) and use the 500 hardest 5-hop problems. (The 3-hop problems used for fine-tuning are described in Section 4).
@@ A.2 Detailed Baseline Methods
Feng et al. (2023) An approach where the language model is trained to directly generate symbolic reasoning steps, emulating a logical solver and bypassing an explicit natural language-to-symbol parsing stage.
@@
Creswell et al. (2022) A method that performs multi-step reasoning through an iterative process, alternating between a "selection" step to identify relevant facts from the context and an "inference" step to derive new conclusions.
@@
Pan et al. (2023) A neuro-symbolic framework where an LLM first translates a problem into a symbolic representation. A deterministic solver then performs the logical inference, and the LLM interprets the result back into natural language. It includes a self-refinement mechanism that uses solver feedback to correct translation errors.
@@
Xu et al. (2024) A framework that integrates symbolic logic into the Chain-of-Thought process. It translates the problem into a symbolic format and then generates reasoning steps using formal inference rules, often employing a verifier to check the logical soundness of the chain.
@@ B.3 Task 1: Correctness Spanning Steps
  • For each test sample, we identify the smallest $i$ such that the classifier correctly predicts the label at step $i$ but fails at step i1i-1.
The correctness spanning length is defined as nin-i, capturing how early the model “knows” the correct answer.
@@ B.5 Task 3: Next-Step Derivability
Appendix C Additional Experiments
Model Setting FLD FOLIO Multi- LogiEval ProntoQA
[…]
Few-shot 44.0 81.8 73.6 99.8
SFT-NL 75.4 64.0 48.7 99.4
SFT-Symb-Struct 58.6 59.1 61.0 99.6
SFT-Symb-Filter 63.6 57.1 64.1 99.8
Qwen-3- 8B SFT-Symb-Direct 54.6 56.2 63.1 75.2
Table 5: Overall benchmark accuracy on Qwen-2.5-7B-Math-Instruct and Qwen-3-8B across training settings. Per-model column-wise maxima are in bold.
Model Setting All Valid All Relevant All Atomic
[…]
SFT-NL 6.5 60.4 2.0
[…]
SFT-Symb-Filter 12.4 3.6 5.6
[…]
SFT-NL 17.3 84.5 3.5
[…]
SFT-Symb-Filter 19.1 2.3 7.6
[…]
Table 6: Stepwise soundness after </think> on Qwen-2.5-7B-Math-Instruct and Qwen-3-8B. Per-model column-wise maxima are in bold.
Model Setting CSS RFI NSD
[…]
SFT-NL 8.6 6.2 42.3
[…]
SFT-Symb-Filter 8.3 8.6 37.1
[…]
SFT-NL 9.8 23.5 36.1
[…]
SFT-Symb-Filter 10.0 14.8 39.2
[…]
Table 7: Representation-level probing (CSS, RFI, NSD) on Qwen-2.5-7B-Math-Instruct and Qwen-3-8B. Per-model column-wise maxima are in bold; the -- row denotes the original model.
This section provides further details and aggregate analysis of our experimental results.
Aggregate Findings Across Models.
To broaden our findings, we evaluated two additional models, Qwen-2.5-Math-7B and Qwen-3-8B. The results, shown in Tables 5, 6 and 7, combined with our primary experiments, yield several key aggregate takeaways:
  • Natural-language SFT is the most reliable path to higher accuracy. Across all four base models and four datasets, SFT-NL consistently delivers the largest and most robust gains in final-answer correctness, reaffirming that aligning with the model’s pre-training modality is most effective for generalization.
  • Symbolic curricula with redundant facts improve chain integrity. SFT settings that retain logically valid but extraneous premises (Symb-Direct and Symb-Struct) tend to improve All-Valid and All-Atomic scores. This suggests that exposing models to "noisy but sound" proof environments encourages more careful and explicit rule application.
  • CSS is more sensitive to architecture and RL than to SFT. The larger, RL-enhanced Qwen-3-8B model achieves a notably higher base CSS score than other models. However, its CSS remains largely unchanged by SFT, providing strong evidence that deeper latent reasoning depends more on model design and training regimes like RL, rather than supervised instruction tuning.
Detailed Analysis of Performance by Reasoning Depth.
A more granular breakdown of performance by reasoning depth, presented in Table 8, reveals further nuances in these trends. While models fine-tuned with natural language supervision (e.g., Llama-3.1-SFT-NL achieving 89.5% accuracy for 0-3 steps on FLD) perform strongly on tasks with shallower reasoning depths, their symbolic reasoning counterparts tend to exhibit greater resilience as the complexity and number of reasoning steps increase. For instance, on FLD problems requiring 16-19 steps, Llama-3.1-SFT-Symb-Filter (62.5%) and Llama-3.1-SFT-Symb-Struct (58.5%) maintain higher accuracy compared to Llama-3.1-SFT-NL (46.0%), highlighting the benefit of symbolic formats for robust long-chain inference.
@@ Appendix C Additional Experiments
[…]
@@ Appendix F Prompt Template
Appendix G Use of AI Assistants in Manuscript Preparation
During the preparation of this manuscript, we utilized large language models, including GPT and Gemini, as writing assistants. Their role was limited to improving the clarity, grammar, and readability of the text. This included tasks such as rephrasing sentences, correcting spelling errors, and ensuring stylistic consistency. The core scientific contributions, experimental design, and analysis presented in this paper are solely the work of the human authors, who take full responsibility for the final content.
@@ Appendix G Use of AI Assistants in Manuscript Preparation
[…]
@@ Conclusion:
Figure 14: Prompt template for evaluating step atomicity. Placeholders: {premises_str} (a string listing the premises), {concl_text_full} (a string representing the conclusion). The model is expected to return ’true’ or ’false’ indicating if the inference from premises to conclusion is a single, indivisible logical step.